# Leveraging LLMs, IDEs, and Semantic Embeddings for Automated Move Method Refactoring

### v1

@@ IV-B Effectiveness of LLMs (RQ1)
Evaluation Metrics. Using these datasets, we evaluated the recommendations made by the vanilla LLM and identified the hallucinations, as defined in Definition III.4: H1 – target class does not exist in the project; H2 – moving the method to the target class is mechanically infeasible; and H3 – violating preconditions in Section III-B.
Experimental Setup. We use the vanilla implementation of GPT-4o, a state-of-the-art LLM from OpenAI [32]. While MM-assist is model agnostic (i.e., we can simply swap different models), we chose GPT-4o because other researchers [39, 40] show that it outperforms other LLMs when used for refactoring tasks. GPT-4o is also widely adopted in many software-engineering tools [41, 42, 43, 44]. We designed our experimental setup to assess the model’s inherent capabilities in understanding and recommending MoveMethod without additional context or task-specific tuning. We formulated a prompt where we provided the source code in a given host class and asked the LLM which methods to move and where. We set the LLM temperature parameter to 0 to obtain deterministic results.
@@
Results. Table I illustrates the distribution of valid suggestions and different types of hallucinations produced by the vanilla LLM for both synthetic and real-world datasets. We observed a prevalence in all three types of hallucinations: H1, H2, and H3 (as defined in Definition III.4) Crucially, actuating any of these hallucinations would lead to broken code, compilation errors, or degraded software design.
@@ IV-C Effectiveness of MM-assist (RQ2)
Baseline Tools. We directly compare with the best-in-class tools: JMove [10] is a state-of-the-art static analysis tool, feTruth [16] is the former best in class tool that uses ML/DL, and HMove [30] is a recently introduced state-of-the-art tool that uses graph neural networks to generate suggestions and LLM to check refactoring preconditions. HMove has been shown to outperform all previous tools. We also compare with the Vanilla-LLM (GPT 4o), which represents the standard LLM solution (without using MM-assist’s enhancements). We went the extra mile to ensure the most fair comparison: we consulted with HMove, feTruth, and JMove authors to ensure the optimal tool’s settings and clarified with the authors when their tools did not produce the expected results. We are grateful for their assistance.
Evaluation Metrics. For evaluation, we employ recall-based metrics, following an approach similar to that used in the evaluation of JMove [10], a well-established tool in this domain. In our setting, recall is a more suitable metric (against precision) because it measures how many relevant recommendations are retrieved, avoiding the need for subjective judgments about whether a recommendations is a false positive. Furthermore, to provide actionable recommendations and avoid overwhelming the developer, we use recall@k, which returns a small number of recommendations (k). Furthermore, recall@k is similar to precision by evaluating the quality of a limited set of recommendations.
@@
[…]
Experimental Setup. We trigger all tools on each host class from the gold set (both synthetic and the real-world dataset). To account for the inherent non-determinism in LLMs, we ran the vanilla LLM with multiple temperature values (0, 0.5, 1) for five runs each. Additionally, we ran MM-assist three times. We report the average and standard deviation for both vanilla LLM and MM-assist. Considering the number of entries in the datasets, given that JMove can take a long time to run (12+ hours on a large project), we cutoff its execution after 1 hour. HMove takes both the method and candidate target class as input, and returns a probability score indicating whether to move the method. As a result, it becomes impractical to recommend moving static methods, because there can be thousands of (method, target-class) pairs – each taking a minute on average to classify. For example, to recommend moving a single static-method in the Elasticsearch project with 21615 target classes would require $\approx 14$ days to process. Thus, we limit HMove to instance move method recommendations.
All the tools generate a ranked list of MoveMethod suggestions. We compared these recommendations against the ground truth to calculate recall for each tool using the evaluation metrics presented earlier – $Recall_{M}$, $Recall_{C}$, and $Recall_{MC}$.
@@
Vanilla-LLM 71$\pm$275$\pm$179$\pm$270$\pm$170$\pm$170$\pm$153$\pm$255$\pm$257$\pm$2%
MM-assist 72$\pm$179$\pm$080$\pm$091$\pm$197$\pm$098$\pm$167$\pm$073$\pm$075$\pm$0%
Results on the Synthetic Corpus. Table II shows the effectiveness of MM-assist and baseline tools on the synthetic dataset. MM-assist demonstrates superior performance across many of the recall metrics compared to other tools, especially $Recall_{MC}@1$ (1.7x to 30x improvement) – the most comprehensive measure assessing both correct method and target class identification. While JMove exhibits high accuracy in target class identification ($Recall_{C}@1$ = 97%), it shows limitations in method identification. HMove demonstrates comparable performance as JMove in identifying misplaced methods (i.e., $Recall_{M}$) but does not match JMove’s ability to recommend the correct target class (i.e., $Recall_{C}$). Consequently, its combined $Recall_{MC}$ is lower than JMove. Interestingly, feTruth achieves perfect $Recall_{C}$ but extremely low $Recall_{M}$ (2% – 3%) despite being very prolific in recommending as many as 67 methods to be moved from a class. When it does correctly identify a method, it accurately suggests the target class. However, because it rarely identifies the misplaced methods themselves, the overall $Recall_{MC}$ remains low. We confirmed this paradox with feTruth authors. Interestingly, the Vanilla-LLM shows comparable performance to MM-assist in method identification ($Recall_{M}$). This could be because the LLMs have been pre-trained on the synthetic dataset, and memorize their responses as a result.
Results on the Real-world Corpus. With the real-world dataset performed by open-source developers, we found a wider difference in performance. First, we distinguish between cases when the MoveMethod target was an instance or static method – shedding light on the effectiveness of MM-assist in different usage scenarios. Second, we distinguish between small and large classes based on their method count. Our analysis of the real-world oracle reveals a heavy-tail distribution – we label classes with fewer than 15 methods (90th percentile across all projects) as Small Classes (avg. 6 methods/class) and the rest as Large Classes (avg. 48 methods/class).
@@ IV-D Runtime performance of MM-assist (RQ3)
Experimental Setup. We used both the synthetic and real-world corpus employed in other RQs to measure the time-taken for each tool to produce recommendations. To understand what components of MM-assist take the most time, we also measured the amount of time it took to generate responses from the LLM, and the time it took to process suggestions. To ensure real-world applicability, we conducted these measurements using the MM-assist plugin for IntelliJ IDEA, mirroring the actual usage scenario for developers. We conducted all experiments on a commodity laptop, an M1 MacBook Air with 16GB of RAM.
Results. Our empirical evaluation demonstrates that MM-assist achieves an average runtime of 27.5 seconds for generating suggestions. The primary computational overhead stems from the LLM API interactions consuming approximately 9 seconds. In our experience with JMove, on the larger projects in our real-world dataset, JMove takes several hours (up to 24 hours) to complete running, thus we imposed the 1-hour cutoff time. Similarly, HMove also takes an average of 80min to execute on a single entry in our dataset – it needs to be triggered on all possible $<$method, target class$>$ pairs (avg. 145 pairs per class). In extreme cases where the host class was large (>10K LOC), HMove took 4 whole days to execute on a single entry in our dataset. Out of the box, feTruth is also slow and can take 12+ hours to run on large projects. With the help of feTruth authors, we were able to run it on a single class at a time – this takes an average 6 minutes per class. Thus, compared with the baselines, MM-assist is two, two, and one orders of magnitude faster than JMove, HMove, and feTruth, respectively. Thus, it is practical.
@@ IV-E Usefulness of MM-assist (RQ4)
Dataset. We made the deliberate choice to have participants use projects with which they were familiar. This decision was grounded in several key considerations. First, familiarity with their codebases enables them to make more informed judgments about the appropriateness and potential impact of the suggested refactorings. Second, using personal projects enhances the validity of our study, as it closely mimics real-world scenarios where developers refactor code they have either authored or maintained extensively. Third, this approach allows us to capture a diverse range of project types, sizes, and domains, potentially uncovering insights that might be missed in a more constrained, standardized dataset.
Experimental Setup. 30 students (25 Master’s and 5 Ph.D. students) volunteered to participate in our study. Based on demographic information provided by the participants, 73% have industrial experience. All participants, with the exception of two, have experience with the Java programming language. Finally, the majority of participants (24 out of 30) have prior experience with refactoring.
We instructed the participants to use MM-assist for a week and run it on at least ten different Java classes from projects they work on. For each class they selected, MM-assist provided up to three MoveMethod recommendations. We chose to present three recommendations to strike a balance between variety and practicality. Afterward, they sent us the fine-grained telemetry data from the plugin usage. For confidentiality reasons, we anonymize the data by stripping away any sensitive information about their code. We collected usage statistics from each invocation of the plugin on each class. In particular, we collected this information: how the users rated each individual recommendation and whether they finally changed their code based on the recommendation.
Participants rated each recommendation on a 6-point Likert scaled ranging from (1) Very unhelpful to (6) Very helpful. We chose this 6-point Likert scale to force a non-neutral stance, encouraging participants to lean towards either a positive or negative assessment. We asked the participants to rate the MM-assist’s recommendations while they were fresh in their minds, right after they analyzed each recommendation.
@@
Results. 30 participants applied MM-assist on 350 classes. We found that, in 290 classes the participants positively rated one of the recommendations (82.8% of the time). Moreover, the users accepted and applied a total of 216 refactoring recommendations on their code, i.e. 7 refactorings per user, on average. This shows that our tool is effective at generating useful recommendations that developers, who are familiar with their code, accept.
@@ V Discussion
External Validity: This concerns the generalization of our results. Because we rely on a specific LLM (GPT-4o), it may impact the broader applicability of our findings. We anticipate that advancements in LLM technology will improve overall performance, though this needs to be verified empirically. Second, MM-assist currently focuses on Java code. Although our approach is conceptually language- and refactoring-agnostic, extending to additional refactoring types and languages requires adapting three key components: (1) static analysis for validating refactoring preconditions, (2) semantic analysis of code relationships, and (3) refactoring execution mechanics. Using protocols like the Language Server Protocol (LSP) [45] can simplify handling language-specific features, facilitating broader applicability. Future work will explore the effectiveness of our tool across various languages and refactorings.
@@ VI Related Work
Most importantly, MM-assist attacks the problem in a drastically different way. Previous tools compute whole project dependencies (which is computationally expensive and doesn’t scale) and then produce a confidence score for each method in the project. Thus, they treat this as a classification, not a recommendation problem: they produce 57 recommendations on average to move out a given class (many of which are unuseful), which puts tremendous analysis burden on the programmer. In contrast, MM-assist offers up to 3 recommendations per class, aligned with how expert developers refactor code.
@@
The prevalence of hallucinations in LLM-based refactoring is widely studied. Pomian et al. [55, 39] investigated hallucinations in ExtractMethod refactoring, while [40] analyzed hallucinations in Python code modifications. These studies consistently show that LLMs can hallucinate during refactoring tasks, substantiating our findings, where LLMs hallucinated in 80% of the cases when suggesting MoveMethod. This highlights the necessity of robust validation mechanisms, which are integral to our MM-assist, ensuring the reliability and safety of the suggestions generated by LLMs.
@@ VII Conclusion
The key to unleashing these breakthroughs is combining static and semantic analysis to (i) eliminate LLM hallucinations and (ii) focus its laser. MM-assist checks refactoring preconditions automatically which cuts down the LLM hallucinations. By leveraging semantic embedding into a RAG approach, MM-assist narrows down the context for the LLM so that it can focus its laser on a small number of high-quality prospects. This was instrumental in picking the right candidate from industrial large scale projects. We hope that these techniques inspire others to solve many other refactoring recommendation domains such as splitting large classes or packages.
@@ Acknowledgements
We are grateful for the constructive feedback from the members of the AI Agents team at JetBrains Research. This research was partially funded through the NSF grants CNS-1941898, CNS-2213763, 2512857, 2512858, the Industry-University Cooperative Research Center on Pervasive Personalized Intelligence, and a gift grant from NEC.

### v2

@@ IV-B Effectiveness of LLMs (RQ1)
IV-B1 Evaluation Metrics
Using these datasets, we evaluated the recommendations made by the vanilla LLM and identified the hallucinations, as defined in Definition˜III.4: H1 – target class does not exist in the project; H2 – moving the method to the target class is mechanically infeasible; and H3 – violating preconditions in Section˜III-B.
IV-B2 Experimental Setup
We use the vanilla implementation of GPT-4o, a state-of-the-art LLM from OpenAI [33]. We used a version that was released on May 13, 2024. While MM-assist is model agnostic (i.e., we can simply swap different models), we chose GPT-4o because researchers [41, 42] show that it outperforms other LLMs when used for refactoring tasks. GPT-4o is also widely adopted in many software-engineering tools [43, 44, 45, 46]. We designed our experimental setup to assess the model’s inherent capabilities in understanding and recommending MoveMethod without additional context or task-specific tuning. We formulated a prompt where we provided the source code in a given host class and asked the LLM which methods to move and where. We set the LLM temperature parameter to 0 to obtain deterministic results.
@@ IV-B2 Experimental Setup
IV-B3 Results
Table˜I illustrates the distribution of valid suggestions and different types of hallucinations produced by the vanilla LLM for both synthetic and real-world datasets. We observed a prevalence in all three types of hallucinations: H1, H2, and H3 (as defined in Definition˜III.4) Crucially, actuating any of these hallucinations would lead to broken code, compilation errors, or degraded software design.
@@ IV-C Effectiveness of MM-assist (RQ2)
IV-C1 Baseline Tools
We directly compare with the best-in-class tools: JMove [10] is a state-of-the-art static analysis tool, feTruth [16] is the former best in class tool that uses ML/DL, and HMove [30] is a recently introduced state-of-the-art tool that uses graph neural networks to generate suggestions and LLM to check refactoring preconditions. HMove has been shown to outperform all previous tools. We also compare with the Vanilla-LLM (GPT 4o), which represents the standard LLM solution (without using MM-assist’s enhancements). We went the extra mile to ensure the most fair comparison: we consulted with HMove, feTruth, and JMove authors to ensure the optimal tool’s settings and clarified with the authors when their tools did not produce the expected results. We are grateful for their assistance. Further, we also replicated the recall numbers presented in each tool’s paper, validating our experimental setup.
IV-C2 Evaluation Metrics
For evaluation, we employ recall-based metrics, following an approach similar to that used in the evaluation of JMove [10], a well-established tool in this domain. In our setting, recall is a more suitable metric (against precision) because it measures how many relevant recommendations are retrieved, avoiding the need for subjective judgments about whether a recommendations is a false positive. Furthermore, to provide actionable recommendations and avoid overwhelming the developer, we use recall@k, which returns a small number of recommendations (k). Furthermore, recall@k is similar to precision by evaluating the quality of a limited set of recommendations.
@@ IV-C2 Evaluation Metrics
For each recall metric, we calculate Recall@k for the top $k$ recommendations, where k $\in$ {1, 2, 3}.
IV-C3 Experimental Setup
We trigger all tools on each host class from the gold set (both synthetic and the real-world dataset). To account for the inherent non-determinism in LLMs, we ran the vanilla LLM with multiple temperature values (0, 0.5, 1) for five runs each. Additionally, we ran MM-assist three times. We report the average and standard deviation for both vanilla LLM and MM-assist. Considering the number of entries in the datasets, given that JMove can take a long time to run (12+ hours on a large project), we cutoff its execution after 1 hour. HMove takes both the method and candidate target class as input, and returns a probability score indicating whether to move the method. As a result, it becomes impractical to recommend moving static methods, because there can be thousands of (method, target-class) pairs – each taking a minute on average to classify. For example, to recommend moving a single static-method in the Elasticsearch project with 21615 target classes would require $\approx 14$ days to process. Thus, we limit HMove to instance move method recommendations. We do not penalize any tool for running out of time – we do not compute recall rates on those data points. For example in Table 3, JMove only finished computing for 24 out of 70 Large Classes, so we only report JMove ’s Recall for those 24 Classes (i.e., we do not compute nor report a zero recall for the 46 Classes it ran out of time).
All the tools generate a ranked list of MoveMethod suggestions. We compared these suggestions against the ground truth to calculate recall for each tool using the evaluation metrics presented earlier – $Recall_{M}$, $Recall_{C}$, and $Recall_{MC}$.
@@ IV-C3 Experimental Setup
[…]
IV-C4 Results on the Synthetic Corpus
Table II shows the effectiveness of MM-assist and baseline tools on the synthetic dataset. MM-assist demonstrates superior performance across many of the recall metrics compared to other tools, especially $Recall_{MC}@1$ (1.7x to 30x improvement) – the most comprehensive measure assessing both correct method and target class identification. While JMove exhibits high accuracy in target class identification ($Recall_{C}@1$ = 97%), it shows limitations in method identification. HMove demonstrates comparable performance as JMove in identifying misplaced methods (i.e., $Recall_{M}$) but does not match JMove’s ability to recommend the correct target class (i.e., $Recall_{C}$). Consequently, its combined $Recall_{MC}$ is lower than JMove. Interestingly, feTruth achieves perfect $Recall_{C}$ but extremely low $Recall_{M}$ (2% – 3%) despite being very prolific in recommending as many as 67 methods to be moved from a class. When it does correctly identify a method, it accurately suggests the target class. However, because it rarely identifies the misplaced methods themselves, the overall $Recall_{MC}$ remains low. We confirmed this paradox with feTruth authors. Interestingly, the Vanilla-LLM shows comparable performance to MM-assist in method identification ($Recall_{M}$). This could be because the LLMs have been pre-trained on the synthetic dataset, and might have memorized their responses – thus becoming necessary to use an LLM uncontaminated dataset (see next).
IV-C5 Results on the Real-world Corpus
With the real-world dataset performed by open-source developers, we found a wider difference in performance. First, we distinguish between cases when the MoveMethod target was an instance or static method – shedding light on the effectiveness of MM-assist in different usage scenarios. Second, we distinguish between small and large classes based on their method count. Our analysis of the real-world oracle reveals a heavy-tail distribution – we label classes with fewer than 15 methods (90th percentile across all projects) as Small Classes (avg. 6 methods/class) and the rest as Large Classes (avg. 48 methods/class).
@@ IV-D Runtime performance of MM-assist (RQ3)
IV-D1 Experimental Setup
We used both the synthetic and real-world corpus employed in other RQs to measure the time taken for each tool to produce recommendations. To understand what components of MM-assist take the most time, we also measured the amount of time it took to generate responses from the LLM, and the time it took to process suggestions. To ensure real-world applicability, we conducted these measurements using the MM-assist plugin for IntelliJ IDEA, mirroring the actual usage scenario for developers. We conducted all experiments on a commodity laptop, an M1 MacBook Air with 16GB of RAM.
IV-D2 Results
Our empirical evaluation demonstrates that MM-assist achieves an average runtime of 27.5 seconds for generating suggestions. The primary computational overhead stems from the LLM API interactions consuming approximately 9 seconds. In our experience with JMove, on the larger projects in our real-world dataset, JMove takes several hours (up to 24 hours) to complete running, thus we imposed the 1-hour cutoff time. Similarly, HMove also takes an average of 80min to execute on a single entry in our dataset – it needs to be triggered on all possible <<method, target class>> pairs (avg. 145 pairs per class). In extreme cases where the host class was large (>10K LOC), HMove took 4 whole days to execute on a single entry in our dataset. Out of the box, feTruth is also slow and can take 12+ hours to run on large projects. With the help of feTruth authors, we were able to run it on a single class at a time – this takes an average 6 minutes per class. Thus, compared with the baselines, MM-assist is two, two, and one order(s) of magnitude faster than JMove, HMove, and feTruth, respectively. Thus, it is practical.
@@ IV-E Usefulness of MM-assist (RQ4)
IV-E1 Dataset
We made the deliberate choice to have participants use projects with which they were familiar. This decision was grounded in several key considerations. First, familiarity with their codebases enables them to make more informed judgments about the appropriateness and potential impact of the suggested refactorings. Second, using personal projects enhances the validity of our study, as it closely mimics real-world scenarios where developers refactor code they have either authored or maintained extensively. Third, this approach allows us to capture a diverse range of project types, sizes, and domains, potentially uncovering insights that might be missed in a more constrained, standardized dataset.
IV-E2 Experimental Setup
30 students (25 Master’s and 5 Ph.D. students) volunteered to participate in our study. Based on demographic information provided by the participants, 73% have industrial experience. All participants, with the exception of two, have experience with the Java programming language. Finally, the majority of participants (24 out of 30) have prior experience with refactoring.
We instructed the participants to use MM-assist for a week and run it on at least ten different Java classes from their projects. The selection of these classes was left to the discretion of the participants, with the guidance to choose files they had either authored or were familiar with. For each class they selected, MM-assist provided up to three MoveMethod recommendations. We chose to present three recommendations to strike a balance between variety and practicality.
The user study involved setting up our IntelliJ plugin, executing our tool on a class to get refactoring suggestions, rating the suggestions live inside the IDE, and lastly, providing feedback by filling out our survey. Afterward, they sent us the fine-grained telemetry data from the plugin usage. For confidentiality reasons, we anonymize the data by stripping away any sensitive information about their code. We collected usage statistics from each invocation of the plugin on each class. In particular, we collected this information: how the users rated each individual recommendation and whether they finally changed their code based on the recommendation.
Participants rated each recommendation on a 6-point Likert scale ranging from (1) Very unhelpful to (6) Very helpful. We chose this 6-point Likert scale to force a non-neutral stance, encouraging participants to lean towards either a positive or negative assessment. We asked the participants to rate the MM-assist’s recommendations while they were fresh in their minds, right after they analyzed each recommendation.
@@ IV-E2 Experimental Setup
IV-E3 Results
30 participants applied MM-assist on 350 classes. MM-assist analyzed 1,143 host classes where developers did not have prior knowledge of refactoring opportunities. MM-assist recommended refactorings in 350 of them, with an average of 1.7 recommendations/class. MM-assist did not deem 793 typical classes to require refactoring, avoiding unnecessary developer effort. We found that, in 290 classes the participants positively rated one of the recommendations (82.8% of the time). Moreover, the users accepted and applied a total of 216 refactoring recommendations on their code (out of 354 total recommendations), i.e. 7 refactorings per user, on average. This shows that our tool is effective at generating useful recommendations that developers, who are familiar with their code, accept.
@@ V Discussion
External Validity: This concerns the generalization of our results. Because we rely on a specific LLM (GPT-4o), it may impact the broader applicability of our findings. We anticipate that advancements in LLM technology will improve overall performance, though this needs to be verified empirically. Second, MM-assist currently focuses on Java code. Although our approach is conceptually language- and refactoring-agnostic, extending to additional refactoring types and languages requires adapting three key components: (1) static analysis for validating refactoring preconditions, (2) semantic analysis of code relationships, and (3) refactoring execution mechanics. Using protocols like the Language Server Protocol (LSP) [47] can simplify handling language-specific features, facilitating broader applicability. Future work will explore the effectiveness of our tool across various languages and refactorings. While we utilize IntelliJ’s refactoring engine for validation, we acknowledge it may contain latent bugs [48]. To mitigate this, MM-assist employs a sanity check to pre-filter invalid move methods. Therefore, the reliability of our approach depends not only on the correctness of IntelliJ’s engine but also on the effectiveness of our pre-filtering stage.
@@ VI Related Work
Most importantly, MM-assist attacks the problem in a different direction. Previous tools compute whole project dependencies (which is computationally expensive and doesn’t scale) and then produce a confidence score for each method in the project. Thus, they treat this as a classification, not a recommendation problem: they produce 57 recommendations on average to move out a given class (many of which are unuseful), which puts tremendous analysis burden on programmers. In contrast, MM-assist offers at most 3 recommendations per class, aligned with how expert developers refactor code.
@@
The prevalence of hallucinations in LLM-based refactoring is widely studied. Pomian et al. [58, 41] investigated hallucinations in ExtractMethod refactoring, while [42] analyzed hallucinations in Python code modifications. These studies consistently show that LLMs can hallucinate during refactoring tasks, substantiating our findings, where LLMs hallucinated in 80% of the cases when suggesting MoveMethod. This highlights the necessity of robust validation mechanisms, which are integral to our MM-assist, ensuring the reliability and safety of the suggestions generated by LLMs. MANTRA [59] introduces a multi-agent approach with RAG to automate refactoring. However, MANTRA uses LLM to generate the refactored code directly, guiding this process with a RAG system that retrieves examples of correct refactorings from open-source projects to serve as few-shot prompts. In contrast, we leverage the LLM only to identify and propose refactoring candidates, while delegating the safe execution of the code transformation to the IDE. Our application of RAG also differs, as we use it to narrow down the search and retrieve a smaller number of potential target classes.
@@ VII Conclusion
The key to unleashing these breakthroughs is combining static and semantic analysis to (i) eliminate LLM hallucinations and (ii) focus its laser. MM-assist checks refactoring preconditions automatically which cuts down the LLM hallucinations. By leveraging semantic embedding into a RAG approach, MM-assist narrows down the context for the LLM so that it can focus on a small number of high-quality prospects. This was instrumental in picking the right candidate from industrial large scale projects. We hope that these techniques inspire others to solve other refactoring recommendation domains, e.g., splitting large classes or packages.
VIII Data Availability
To ensure the verifiability of our work and to provide a foundation on which others in the community can build upon, we have made our tool, MM-assist, publicly available in our replication package  [31]. This includes the complete source code, the IntelliJ plugin implementation, comprehensive documentation detailing the setup and usage instructions, a demo of MM-assist in action, the exact LLM prompts that we use, and all datasets we use in our evaluation.
@@ Acknowledgements
We are grateful for the constructive feedback from the members of the AI Agents team at JetBrains Research and the anonymous conference reviewers. This research was partially funded through the US National Science Foundation (NSF) grants CNS-1941898, CNS-2213763, 2512857, 2512858, the Industry-University Cooperative Research Center on Pervasive Personalized Intelligence (PPI), and a gift grant from NEC. Tien N. Nguyen was supported in part by the NSF grant CNS-2120386, and the National Security Agency (NSA) grant NCAE-C-002-2021.